### v1

@@ Appendix B Proof of Proposition 4
$(1-a)T-b\leq H_{v}\leq(1+a)T+b$ (42)
where $T=-1/2\Delta$ is the kinetic energy operator. Moreover, the spectrum of the Hamiltonian $H_{v}$ (and also of $T$) contains only discrete values due to the compact embedding $Q(H_{v})=H^{1}(\mathbb{T}^{N})\hookrightarrow L^{2}(\mathbb{T}^{N})$. Then we can express $H_{v}$ and $T$ in spectral decomposition
@@
where we additionally require $\lambda_{1}\leq\lambda_{2}\leq\ldots$ and $\mu_{1}\leq\mu_{2}\leq\ldots$. Then, we get the following lemma.
@@
The $n$’th eigenvalue of $H_{v}$ has the lower and upper bound, $(1-a)\mu_{n}-b\leq\lambda_{n}\leq(1+a)\mu_{n}+b$.
@@
Let $\phi_{j}$ be the states introduced in Eq. (43). By the min-max-principle for self-adjoint operators and by Eq. (42) we get
@@
For the lower bound we proceed in the same way but we use the max-min-principle instead to get
@@
We use the lower and upper bound of the eigenvalues to proof that the kinetic energy expectation is finite.
@@
For all, $v\in H^{-1}$ we have $\mathrm{Tr}\{T\Gamma_{v}\}<\infty$.
@@
The Gibbs state is given by $|\Gamma_{v}=\frac{1}{Z(v)}\sum_{j=1}^{\infty}\mathrm{e}^{-\beta\lambda_{j}}\ket{\psi_{j}}\bra{\psi_{j}}$. We first show that the energy expectation is finite by using Lemma 19. That is,
-$\infty$ $<\mathrm{Tr}\{H_{v}\Gamma_{v}\}=\frac{1}{Z(v)}\sum\limits_{j=1}^{\infty}\lambda_{j}\mathrm{e}^{-\beta\lambda_{j}}\leq\frac{1}{Z(v)}\sum\limits_{j=1}^{\infty}\big{[}(1+a)\mu_{j}+b\big{]}\mathrm{e}^{-\beta[(1-a)\mu_{j}-b]}$
$=\frac{(1+a)}{Z(v)}\mathrm{e}^{\beta b}\sum\limits_{j=1}^{\infty}\mu_{j}\mathrm{e}^{-\beta(1-a)\mu_{j}}+\frac{\mathrm{e}^{\beta b}b}{Z(v)}\sum\limits_{j=1}^{\infty}\mathrm{e}^{-\beta(1-a)\mu_{j}}.$
A similar computation as in Eq. (38) shows that the two sums are finite and we get $\mathrm{Tr}\{H_{v}\Gamma_{v}\}<\infty$. Using Eq. (42) and positivity of the kinetic energy operator gives
$-b\leq(1-a)\mathrm{Tr}\{T\Gamma_{v}\}-b\leq\mathrm{Tr}\{H_{v}\Gamma_{v}\}<\infty,$
@@
For simplicity we introduce $\alpha_{j}=\mathrm{e}^{-\beta\lambda_{j}}/Z(v)$ such that the Gibbs state is $|.\Gamma_{v}=\sum_{j=1}^{\infty}\alpha_{j}\ket{\psi_{j}}\bra{\psi_{j}}$. Then, its density is given by $\rho=\sum_{j=1}^{\infty}\alpha_{j}\rho_{j}$ with $\psi_{j}\mapsto\rho_{j}$. By Theorem 2 all the pure state densities $\rho_{j}$ are contained in $\mathscr{I}$. We define a sequence $\rho^{n}=\sum_{j=1}^{n}\alpha_{j}\rho_{j}$. Clearly, $\rho^{n}\to\rho$ in $L^{1}(\mathbb{T})$ and $\|\sqrt{\rho^{n}}\rVert_{2}^{2}\leq N$. Since $\int(\nabla\sqrt{\rho})^{2}\mathrm{d}x$ is convex on $\mathscr{I}$ [6, remark after Eq. (1.10)], we find by Jensen’s inequality and by Theorem 2
$\int(\nabla\sqrt{\rho^{n}})^{2}\mathrm{d}x\leq\sum\limits_{j=1}^{n}\alpha_{j}\int(\nabla\sqrt{\rho_{j}})^{2}\mathrm{d}x\leq\sum\limits_{j=1}^{n}\alpha_{j}2T(\psi_{j})\leq 2\mathrm{Tr}\{T\Gamma_{v}\}<\infty,$
where we get a uniform bound by Lemma 20. In particular, $\sqrt{\rho^{n}}$ is a bounded sequence in $H^{1}(\mathbb{T})$ and by the Banach–Alaoglu theorem, there is a subsequence (also denoted by $\sqrt{\rho^{n}}$) and $f\in H^{1}(\mathbb{T})$ such that $\sqrt{\rho^{n}}\rightharpoonup f$ weakly in $H^{1}$. But since $H^{1}(\mathbb{T})$ is compactly embedded in $L^{2}(\mathbb{T})$ there is also a subsequence (again denoted by $\sqrt{\rho^{n}}$) such that $\sqrt{\rho^{n}}\to f$ strongly in $L^{2}(\mathbb{T})$. Then, we have
$\|\rho^{n}-f^{2}\rVert_{1}=\int(\sqrt{\rho^{n}}-f)(\sqrt{\rho^{n}}+f)\mathrm{d}x\leq\|\sqrt{\rho^{n}}-f\rVert_{2}\|\sqrt{\rho^{n}}+f\rVert_{2}\to 0$
since $\|\sqrt{\rho^{n}}+f\rVert_{2}$ is bounded. This means that $\rho^{n}\to f^{2}$ strongly in $L^{1}$ and we conclude $\sqrt{\rho}=f\in H^{1}(\mathbb{T})$. ∎

### v2

@@ Appendix B Proof of Proposition 4
$(1-a)T-b\leq H_{v}\leq(1+a)T+b,$ (49)
where $T=-\frac{1}{2}\Delta$ is the kinetic energy operator. Moreover, the Hamiltonian $H_{v}$ (and also $\Delta/2$) has discrete spectrum due to the compact embedding $Q(H_{v})=\bigl(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr)^{\wedge N}\hookrightarrow\bigl(L^{2}_{\mathbb{C}}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr)^{\wedge N}$. We can express $H_{v}$ and $T$ in their respective spectral decomposition
@@
where we additionally require $\lambda_{1}\leq\lambda_{2}\leq\ldots$ and $\mu_{1}\leq\mu_{2}\leq\ldots$. Then we get the following lemma.
@@
The $n$’th eigenvalue of $H_{v}$ has the following lower and upper bound, $(1-a)\mu_{n}-b\leq\lambda_{n}\leq(1+a)\mu_{n}+b$.
@@
Let $\phi_{j}$ be the states introduced in Eq. (50). By the min-max principle for self-adjoint operators and by Eq. (49) we get
@@
For the lower bound we proceed in the same way but we use the max-min principle instead to get
@@
This gives the desired bounds. ∎
We use the lower and upper bound of the eigenvalues to prove that expectation value of the kinetic energy operator is finite.
@@
For all $v\in H^{-1}(\mathbb{T})$ we have $0\leq\operatorname{Tr}\{T\Gamma_{v}\}<\infty$ and $\infty<\operatorname{Tr}\{H_{v}\Gamma_{v}\}<\infty$.
@@
The Gibbs state is given by $|\Gamma_{v}=\frac{1}{Z(v)}\sum_{j=1}^{\infty}\mathrm{e}^{-\beta\lambda_{j}}\ket{\psi_{j}}\bra{\psi_{j}}$. We first show that the expectation value is finite by using Lemma 15. That is,
-$\infty$ $<\operatorname{Tr}\{H_{v}\Gamma_{v}\}=\frac{1}{Z(v)}\sum\limits_{j=1}^{\infty}\lambda_{j}\mathrm{e}^{-\beta\lambda_{j}}\leq\frac{1}{Z(v)}\sum\limits_{j=1}^{\infty}\big[(1+a)\mu_{j}+b\big]\mathrm{e}^{-\beta[(1-a)\mu_{j}-b]}$ (53)
$=\frac{\mathrm{e}^{\beta b}(1+a)}{Z(v)}\sum\limits_{j=1}^{\infty}\mu_{j}\mathrm{e}^{-\beta(1-a)\mu_{j}}+\frac{\mathrm{e}^{\beta b}b}{Z(v)}\sum\limits_{j=1}^{\infty}\mathrm{e}^{-\beta(1-a)\mu_{j}}.$
A similar computation as in Eq. (44) shows that the two sums are finite and we get $\operatorname{Tr}\{H_{v}\Gamma_{v}\}<\infty$. Using Eq. (49) and positivity of the kinetic energy operator gives
$-b\leq(1-a)\operatorname{Tr}\{T\Gamma_{v}\}-b\leq\operatorname{Tr}\{H_{v}\Gamma_{v}\}<\infty,$ (54)
@@
For simplicity we introduce $\alpha_{j}=\mathrm{e}^{-\beta\lambda_{j}}/Z(v)$ such that the Gibbs state is $|.\Gamma_{v}=\sum_{j=1}^{\infty}\alpha_{j}\ket{\psi_{j}}\bra{\psi_{j}}$. Then, its density is given by $\rho=\sum_{j=1}^{\infty}\alpha_{j}\rho_{j}$ with $\psi_{j}\mapsto\rho_{j}$. By Theorem 2 all the pure state densities $\rho_{j}$ are contained in $\mathscr{I}$. We define a sequence $\rho^{n}=\sum_{j=1}^{n}\alpha_{j}\rho_{j}$. Clearly, $\rho^{n}\to\rho$ in $L^{1}(\mathbb{T})$ and $\|\sqrt{\rho^{n}}\rVert_{2}^{2}\leq N$. Since $\int(\nabla\sqrt{\rho})^{2}\mathrm{d}x$ is convex on $\mathscr{I}$ [28, remark after Eq. (1.10)], we find by Jensen’s inequality and by Theorem 2 that
$\int(\nabla\sqrt{\rho^{n}})^{2}\mathrm{d}x\leq\sum\limits_{j=1}^{n}\alpha_{j}\int(\nabla\sqrt{\rho_{j}})^{2}\mathrm{d}x\leq\sum\limits_{j=1}^{n}\alpha_{j}2T(\psi_{j})\leq 2\operatorname{Tr}\{T\Gamma_{v}\}<\infty,$ (55)
where we get a uniform bound by Lemma 16. In particular, $\sqrt{\rho^{n}}$ is a bounded sequence in $H^{1}(\mathbb{T})$ and by the Banach–Alaoglu theorem there is a subsequence (also denoted by $\sqrt{\rho^{n}}$) and $f\in H^{1}(\mathbb{T})$ such that $\sqrt{\rho^{n}}\rightharpoonup f$ weakly in $H^{1}(\mathbb{T})$. But since $H^{1}(\mathbb{T})$ is compactly embedded in $L^{2}(\mathbb{T})$ there is also a subsequence (again denoted by $\sqrt{\rho^{n}}$) such that $\sqrt{\rho^{n}}\to f$ strongly in $L^{2}(\mathbb{T})$. Then, we have
$\|\rho^{n}-f^{2}\rVert_{1}=\int(\sqrt{\rho^{n}}-f)(\sqrt{\rho^{n}}+f)\mathrm{d}x\leq\|\sqrt{\rho^{n}}-f\rVert_{2}\|\sqrt{\rho^{n}}+f\rVert_{2}\to 0$ (56)
since $\|\sqrt{\rho^{n}}+f\rVert_{2}$ is bounded. This means that $\rho^{n}\to f^{2}$ strongly in $L^{1}(\mathbb{T})$ and we conclude $\sqrt{\rho}=f\in H^{1}(\mathbb{T})$. ∎